# IW-Bench: Evaluating Large Multimodal Models for Converting Image-to-Web

### v1

@@ C.3 Chain-of-Thought
We have witnessed the remarkable capabilities of Large Language Models (LLMs), with their reasoning abilities significantly enhanced through approaches such as Chain-of-Thought (CoT) (Wei et al., 2022), Program-of-Thought (PoT) (Chen et al., 2022), and Inductive Reasoning (Wang et al., 2023a; Tan and Motani, 2023; Guo et al., 2023). For multimodal CoT, recent work includes MCoT Zhang et al. (2023d), HoT Yao et al. (2023), CoCoT Zhang et al. (2024). In this paper, we propose a five-hop multimodal Chain-of-Thought for evaluating multimodal large models on image-to-web domain while we compare other multimodal baselines.
@@ D.1 General prompt for JavaScript Events
Content Unified JavaScript Event: sonclick, onload, onmouseover, onmouseout, onchange, onsubmit, onmousemove, onmouseup, onmousedown, ondblclick, onkeydown, onkeyup, onkeypress, onsubmit, onfocus, onblur, oninput, onload, onresize, onscroll, onunload, ontouchstart, ontouchmove, ontouchend, onerror, oncontextmenu.
@@ Appendix G Web Page Quality and User Experience Questionnaire
Questionnaire 1. Content Comprehensibility 1.1. The content on the web page is easy to understand. 1.2. The information provided is relevant and useful. 1.3. The language used is clear and concise. 2. Layout and Structure 2.1. The layout of the web page is visually appealing. 2.2. The structure of the web page is logical and easy to follow. 2.3. The web page is well-organized. 3. Interactivity and Functionality 3.1. The interactive elements on the web page work as expected. 3.2. The web page is responsive and loads quickly. 3.3. The navigation is intuitive and user-friendly. 4. Overall Satisfaction 4.1. I am satisfied with my overall experience on the web page. 4.2. I would recommend this web page to others. 4.3. I am likely to return to this web page in the future. Rating Scale 1 - Very dissatisfied/Very difficult to achieve 2 - Dissatisfied/Difficult to achieve 3 - Neutral 4 - Satisfied/Easy to achieve 5 - Very satisfied/Very easy to achieve Thank You Thank you for taking the time to complete this questionnaire. Your feedback is valuable and will help us improve the quality of our web pages and user experience.

### v2

@@ C.3 Chain-of-Thought
We have witnessed the remarkable capabilities of Large Language Models (LLMs), with their reasoning abilities significantly enhanced through approaches such as Chain-of-Thought (CoT) (wei2022chain), Program-of-Thought (PoT) (chen2022program), and Inductive Reasoning (wang2023hypothesis; tan2023large). For multimodal CoT, recent work includes MCoT mmcot, HoT hot, CoCoT cocot. In this paper, we propose a five-hop multimodal Chain-of-Thought for evaluating multimodal large models on image-to-web domain while we compare other multimodal baselines.
@@ D.1 General prompt for JavaScript Events
Content Unified JavaScript Event: onclick, onload, onmouseover, onmouseout, onchange, onsubmit, onmousemove, onmouseup, onmousedown, ondblclick, onkeydown, onkeyup, onkeypress, onsubmit, onfocus, onblur, oninput, onload, onresize, onscroll, onunload, ontouchstart, ontouchmove, ontouchend, onerror, oncontextmenu.
@@ Appendix G Web Page Quality and User Experience Questionnaire
Questionnaire 1. Content Comprehensibility $1.$ The content on the web page is easy to understand. 1.2. The information provided is relevant and useful. 1.3. The language used is clear and concise. 2. Layout and Structure 2.1. The layout of the web page is visually appealing. $2.$ The structure of the web page is logical and easy to follow. 2.3. The web page is well-organized. 3. Interactivity and Functionality 3.1. The interactive elements on the web page work as expected. 3.2. The web page is responsive and loads quickly. $3.$ The navigation is intuitive and user-friendly. 4. Overall Satisfaction 4.1. I am satisfied with my overall experience on the web page. 4.2. I would recommend this web page to others. 4.3. I am likely to return to this web page in the future. Rating Scale 1 - Very dissatisfied/Very difficult to achieve 2 - Dissatisfied/Difficult to achieve 3 - Neutral 4 - Satisfied/Easy to achieve 5 - Very satisfied/Very easy to achieve Thank You Thank you for taking the time to complete this questionnaire. Your feedback is valuable and will help us improve the quality of our web pages and user experience.